OpenStudy (anonymous):

PLEASE HELP! SO CONFUSED! Use synthetic substitution to evaluate P(x) = x3 + 3x2 - 6 for x = -1. P(-1) =

OpenStudy (anonymous):

for this one it is much easier to compute directly then by synthetic division

OpenStudy (anonymous):

\[P(-1)=(-1)^3+3(-1)^2-6=-1+3-6=-4\]

OpenStudy (anonymous):

What is p was equal to 5?